Log in
wiki.ucar.edu
Spaces
Hit enter to search
Help
Online Help
Keyboard Shortcuts
Feed Builder
What’s new
About Confluence
Log in
CABL Educational Field Project - EOL
Edit space details
Pages
Blog
Blog
2015
July
June
May
April
Tirga test results
No rad data
GPS at 200m out
Even more TRH 300m problems
TRH hiccups
ehs removed today
/scr/isfs down
Sonic data on 300m tower 4/17
Status update Friday
ehs data turned off
New bug in R code causing problems
Status update Monday
Status update Saturday
Status update Thursday
TRH.300m restarted
Distinct mast effects on SE sonics
Status update Tuesday
Status update Saturday/Sunday
Status update Friday
Picklefork revived
Post-Mortem issues
300m TRH restarted 18:20 MDT
Status update Wednesday
100m cannot keep time
March
February
Browse pages
Configure
Space tools
View Page
A
t
tachments (0)
Page History
Page Information
View Source
Export to PDF
Hide Inline Comments
Blog
2015
April
02
Picklefork revived
Attachments
There are currently no attachments on this page.
Overview
Content Tools
{"serverDuration": 43, "requestCorrelationId": "a5b974812e696d68"}